OpenStudy (anonymous):

What is the remainder when (2x4 – 3x3 + x2 – 3x – 3) ÷ (x – 2) ? 6 3 –9 –66

OpenStudy (anonymous):

\[2x^4 – 3x^3 + x^2 – 3x – 3=2x^4-4x^3+x^3-2x^2+3x^2-6x+3x-6+3\\=(x-2)(2x^3+x^2-3x+3)+3\]

OpenStudy (anonymous):

b is right